# ADAPT: A Pseudo-labeling Approach to Combat Concept Drift in Malware Detection

### v1

@@ X Limitations & Future Work
Evasion and Poisoning Attacks. The effectiveness of our approach may be compromised by adversarial samples designed to evade malware detection classifiers [34]. While we do not explicitly address adversarial examples, integrating adversarial robustness techniques [14] could enhance resilience. Additionally, poisoning attacks can impact continual learning updates [61]; if an adversary crafts a misclassified sample with high confidence, it may increase false negatives or false positives through pseudo-labeling errors. However, semi-supervised learning methods may offer some resistance to label noise [69]. We leave the exploration of adversarial robustness in this context as future work.
Computational Efficiency. ADAPT introduces no additional latency during deployment, as its inference time remains identical to that of the baseline model. However, the training time increases linearly due to data augmentation and mixup strategies. We provide a detailed analysis of the computational overhead in Appendix I.
@@
Catastrophic forgetting. Catastrophic forgetting refers to the tendency of a learned model to lose previously acquired knowledge when trained on new data—a challenge particularly relevant in malware detection [53]. We analyze its impact on the Android malware detection datasets in Appendix H. While our approach does not directly address catastrophic forgetting, future work could jointly integrate mechanisms to handle concept drift and mitigate forgetting. One promising direction is using replay-based methods from continual learning [54]. In our context, this could involve maintaining a replay buffer of class-balanced, confidently pseudo-labeled samples from intermediate time steps, which can be reused during subsequent model fine-tuning.
@@ XI Conclusion
We propose a pseudo-labeling-based approach to mitigate concept drift in malware detection. Our model-agnostic method demonstrates effectiveness across five different datasets, each with varying degrees of distributional shift. Through these experiments, we show how our approach addresses limitations found in prior pseudo-labeling methods applied to malware detection. Additionally, our method is compatible with active learning, offering a pathway for further improvement in handling concept drift.
Appendix A Code Availability
Our work leverages publicly available datasets for malware detection, and we provide detailed explanations of how these datasets were adapted for our experimental setup to ensure reproducibility. Code is available at https://anonymous.4open.science/r/adapt-3006. Dataset splits, experiment logs, and trained models will be made publicly available upon publication to support further research and replication of our results.
Appendix B Self-Training Baselines
DroidEvolver [70] and its improved version, DroidEvolver++ (DE++) [38], are notable prior works in Android malware detection using self-training. However, they suffer from performance degradation due to self-poisoning and have not been thoroughly evaluated across different datasets. Beyond these, there is limited research exploring self-training for concept drift adaptation in the context of malware detection. Therefore, besides DE++, we repurpose and implement three well-known self-training algorithms to evaluate alongside our method.
@@ Appendix B Self-Training Baselines
    Adaptive Random Forest (ARF) [32]: ARF is an online learning algorithm that adapts to concept drift through an ensemble of decision trees. While ARF typically requires access to ground truth labels for drift detection and model updates, we adapt it by introducing a threshold parameter to select pseudo-labeled samples. This threshold is tuned alongside other tree hyperparameters, enabling ARF to operate in a self-training setting.
@@
    DroidEvolver++ (DE++) [38]: DE++ is an enhanced version of DroidEvolver [70], designed to adapt malware detection models to concept drift without continuous ground truth annotations. DE++ employs an ensemble of five linear models—Passive-Aggressive (PA1), Online Gradient Descent (OGD), Adaptive Regularization of Weights (AROW), Regularized Dual Averaging (RDA), and Adaptive Fobos (Ada-Fobos). It uses the ensemble’s predictions to generate pseudo-labels and maintains a fixed-size buffer of recent samples (the app buffer) to monitor model aging. A model is considered “aging” if the proportion of consistent predictions within the buffer falls below a threshold. If this happens, the model is updated using the ensemble’s pseudo-labels.
@@
    Insomnia [10]: Insomnia is a concept drift adaptation method for network intrusion detection, employing co-training with a Nearest Centroid (NC) classifier and an MLP. The NC classifier acts as an oracle, providing pseudo-labels to the MLP for its least confident predictions, which are then used to update the MLP. Both the NC classifier and the MLP are updated using pseudo-labeled data. The authors argue that these two classifiers complement each other: the MLP produces uncertainty estimates, while the NC classifier provides pseudo-labels. However, we observed that the NC classifier performed poorly on the more complex task of malware detection, so we replaced it with an XGBoost model while retaining the MLP. Additionally, we selected the most confident samples rather than the least confident ones, as the latter led to self-poisoning early in the training phase.
@@
    MORSE [69]: MORSE is the state-of-the-art semi-supervised learning method for malware family classification under noisy labels. Since MORSE also employs semi-supervised techniques, we include it as a baseline. It implements the FixMatch [59] algorithm for malware classification, incorporating a novel augmentation strategy that ensures consistency between a weakly augmented pseudo-labeled sample and a strongly augmented version of the same sample.
The first two baseline methods, ARF and DE++, are online methods that process one sample at a time and do not require access to the entire training dataset after the initial training phase. In contrast, the latter two methods—Insomnia and MORSE—use batch training, combining original training data with pseudo-labeled data for model updates. While we primarily assume that training data is available, we also explore scenarios where access to the original training data is restricted, as discussed in Appendix G.
Appendix C Multiclass Classification
C-A Dataset Statistics
Table XIII for the BODMAS dataset and Table XIV for the EMBER dataset present the class-wise statistics for the training, validation, and test sets.
TABLE XIII: Statistics of the BODMAS multiclass classification dataset
ID Family Train Samples Validation Samples Test Samples
@@ 10 Berbew 284 195 1195
In the BODMAS dataset, we use all available samples from the training split for each malware family and randomly select 1,000 samples per month from the Benign class. The same sampling strategy is applied to the validation and test splits. Notably, some malware families, such as Gepys, are relatively rare in the test set, whereas others, such as Berbew, appear more frequently than in the training data.
TABLE XIV: Statics of the EMBER multiclass classification dataset
ID Family Train Samples Test Samples
@@ 10 Tiggre 405 1284
Due to the large dataset size, we cap the maximum number of samples per class at 500 for the EMBER training dataset, including the Benign class. This limitation is also enforced for each month in the test set, ensuring that no class exceeds 500 samples per month. While the training data is approximately balanced across classes, the test data exhibits class imbalance.
C-B Experiments Excluding Benign Class
In practice, malware analysis is often a two-step process. The first step involves binary classification to determine whether a sample is benign or malicious. If classified as malware, the next step is to provide additional context by identifying its malware family [53]. We consider this setting for malware family classification here without including the benign class. We remove the benign samples from the BODMAS (Table XIII) and EMBER datasets(Table XIV) , resulting in a 10-class classification task.
In this setting, we use a single threshold parameter, $\tau_{m}$, in ADAPT, shared across all malware families. Aside from this threshold adjustment, the rest of the algorithm remains unchanged between binary and multiclass classification. This modification reduces the number of hyperparameters in ADAPT from five to four.
TABLE XV: Performance on multiclass classification (10-class) on BODMAS and EMBER datasets.
@@ C-B Experiments Excluding Benign Class
Table XV presents the results of ADAPT and other self-training approaches across both datasets. ADAPT outperforms the baseline XGBoost methods on both datasets, while other self-training methods perform worse. The improvement is especially notable on the EMBER dataset, where ADAPT achieves a 6.5% increase in the F1-score.
TABLE XVI: Performance on multiclass classification (10-class) on BODMAS and EMBER datasets with active learning.
@@
Table XVI presents the results for active learning with a monthly annotation budget of 50 samples. Active learning with uncertainty sampling outperforms the baseline XGBoost model trained with the same annotation budget using random sampling. Additionally, incorporating ADAPT further enhances performance on both datasets. Specifically, integrating ADAPT improves the F1-score by 2.3% on the BODMAS dataset and 6.1% on the EMBER dataset, demonstrating the effectiveness of leveraging unlabeled samples even in the presence of active learning.
Appendix D Support Vector Machine(SVM) Experiments
We conduct additional experiments using SVM baseline models on two Android datasets, APIGraph and Drebin, using the same hyperparameter selection strategy from Section VI-C. The results, including those with ADAPT, are summarized in Table XVII.
The baseline SVM model performs competitively with the baseline XGBoost model on the APIGraph dataset. However, its performance is significantly lower on the Drebin dataset than the XGBoosr baseline.
@@ Appendix D Support Vector Machine(SVM) Experiments
TABLE XVII: Performance comparison on APIGraph and Drebin datasets using SVM model.
Method F1 FPR FNR
APIGraph
SVM 66.3 ± 0.00 1.41 ± 0.00 39.5 ± 0.00
SVM+ADAPT 73.6 ± 0.13 1.47 ± 0.01 29.3 ± 0.08
Drebin
SVM 33.8 ± 0.00 0.70 ± 0.00 76.7 ± 0.00
SVM+ADAPT 48.0 ± 0.54 0.84 ± 0.06 61.5 ± 0.60
Appendix E Statistical Significance
Table XVIII presents the Wilcoxon signed-rank test p-values for comparing different machine learning models (Random Forest, MLP, and XGBoost) across multiple malware datasets. The Wilcoxon test [68] is a non-parametric statistical test used to assess whether two paired samples come from the same distribution. In this case, it evaluates whether the adapted model’s F1-scores significantly differ from the baseline in the test months. Cells are shaded in gray if the p-value is less than 0.05, indicating statistically significant differences. Notably, we observe consistently low p-values for the XGBoost models across all datasets, suggesting a significant effect of adaptation on their performance.
TABLE XVIII: Wilcoxon p-values for different models across datasets. Cells with $p<0.05$ are shaded.
Dataset Metric Random Forest MLP XGBoost
p-values
Drebin p-value $7.63\times 10^{-6}$ $5.34\times 10^{-5}$ $1.53\times 10^{-5}$
APIGraph p-value $1.64\times 10^{-12}$ $1.64\times 10^{-12}$ $1.23\times 10^{-10}$
BODMAS p-value $0.6875$ $0.03125$ $0.03125$
EMBER p-value $0.7646$ $0.03223$ $0.00195$
PDF p-value $0.4375$ $0.15625$ $0.03125$
Table XIX presents the p-values for active learning using ADAPT with the XGBoost baseline model with 50 randomly annotated samples per month.
TABLE XIX: Wilcoxon p-values for active learning with ADAPT vs baseline XGBoost model (50 annotation budget per month). Cells with $p<0.05$ are shaded.
Dataset p-value
Drebin $7.63\times 10^{-6}$
APIGraph $1.64\times 10^{-12}$
Figure 10: Performance on the reduced BODMAS dataset (with eight malware families in the training and validation sets) evaluated on validation and test data: F1 score (left), FPR (middle), and FNR (right).
Appendix F Drifted BODMAS dataset
Since the BODMAS dataset does not exhibit significant drift, we curated an additional dataset to evaluate model performance under more severe drift. For this purpose, we utilized the available malware family information in the dataset. Specifically, we included only the top 8 most frequent malware families from the first three months in the training and validation sets. This corresponds to 5,856 samples or 52.8% of the total malware samples in the training set. We then randomly sampled an equal number of benign samples to include in the training set, as the original dataset has roughly equal distributions of benign and malware samples. The same process was followed for the validation set, comprising the next three months. For the test months, we used the original samples for those months. As a result, the dataset exhibits significant drift starting in the first test month compared to the training and validation data.
We conducted a separate hyperparameter search for this setting and plotted the results of four models in Figure 10. As expected, the baseline XGBoost model deteriorates rapidly starting from month 4 (the first test month) due to a high number of false negatives. In contrast, our proposed method is more resistant to the drift, maintaining relatively high performance, except in month 7, where all models suffer a significant drop in performance.
Appendix G Source-Free Adaptation
@@
In some scenarios, access to the original training data may be unavailable after model training due to privacy concerns, proprietary restrictions (for benign applications), storage constraints, or computational limitations [70]. In such cases, the model needs to be updated using only pseudo-labeled data. We explore this possibility for the XGBoost model on the Drebin and APIGraph datasets. The only modification to our proposed Algorithm 1 is in line 7, where $D_{m}=D_{p}$, i.e., we only have access to pseudo-labeled data after the initial training period. We perform separate hyperparameter optimization for this setting using the same search space.
Figure 11 shows the results on these datasets, comparing the baseline XGBoost model, the model updated in the semi-supervised learning setup with ADAPT, and the source-free adaptation. For the APIGraph dataset, we plot results for the first 18 months of the test set, as performance deteriorates in the long run (with an overall average F1 score of 61.24%). However, the method shows noticeable benefits within this shorter time frame, even in the source-free setting.
For the Drebin dataset, the source-free adaptation improves the average F1 score by 11.3% over the baseline model in the first year, trailing the semi-supervised variant by only 1.76%. On the APIGraph dataset, it improves over the baseline by 0.82% but trails the semi-supervised baseline by 3.97%. These results suggest that ADAPT can mitigate concept drift to some extent, even when the original training data is unavailable. This makes it a viable option for continual learning settings with storage constraints on the training data, and it can be explored in future research [53].
@@
Figure 12: Impact of catastrophic forgetting on the Drebin dataset for XGBoost and MLP models. The performance is shown over the first 12 months of the test dataset. $M_{T}$ represents the model trained in the final month, while $M_{i-1}$ denotes the model trained during the preceding month in the retraining process.
@@
Appendix H Catastrophic Forgetting
Catastrophic forgetting is a significant concern for any continual learning method, where a model may lose previously learned knowledge when trained on new data [53]. Although addressing catastrophic forgetting is beyond the scope of this work, we assess its impact on the Android malware datasets using XGBoost and MLP.
To assess this, we take the final model trained on the last month ($M_{T}$) and evaluate its performance on the first year of test data. We then compare the F1 score, False Positive Rate (FPR), and False Negative Rate (FNR) of the final model against the performance of the model trained during the monthly retraining phase.
@@ Appendix H Catastrophic Forgetting
TABLE XX: Training time (in seconds) of different baseline models with ADAPT
Dataset RF RF + ADAPT MLP MLP + ADAPT XGBoost XGBoost + ADAPT
@@
Appendix I Computational Efficiency
The extent of this increase depends on the dataset size, feature dimensionality, and model complexity. Table XX reports the training time overhead introduced by ADAPT across different models and datasets. The most significant increase is observed for the XGBoost model on the Drebin dataset, where the high feature dimensionality (16,978) substantially impacts computation time. For other datasets, training time typically increases by three to five relative to the baseline model.
Appendix J Data Augmentation & Consistency
We do not explicitly ensure that augmented feature vectors correspond to valid samples, as our augmentation operates in feature space rather than generating samples from a known distribution. Ensuring validity would require modeling complex feature dependencies, which are intractable to the best of our knowledge.
However, we analyze feature correlation structures to assess whether augmentation disrupts dependencies. Specifically, we compute the mean absolute correlation of features in the original dataset to check for strong dependencies. Additionally, we measure the Spearman correlation between the upper triangular elements of the original and augmented correlation matrices to quantify how well feature relationships are preserved.
TABLE XXI: Comparison of Feature Correlations Before and After Augmentation
Dataset Mean $|\mathit{\rho}(X)|$ Spearman ($\mathit{\rho}(X)$, $\mathit{\rho}(X_{\text{aug}})$)
APIGraph 0.0238 0.9247
Drebin 0.0269 0.9911
Bodmas 0.0276 0.9491
PDF 0.0857 0.9975
Table XXI shows feature correlations are generally low across datasets, suggesting weak dependencies. Furthermore, the high Spearman correlation indicates that augmentation does not significantly alter feature relationships. While this does not guarantee individual sample validity, it provides evidence that the overall statistical structure remains consistent.
(a) APIGraph
(b) BODMAS
Figure 14: Validation set performance with and without label consistency as the augmentation mask rate increases.
One way to introduce potentially inconsistent samples is by increasing the strength of augmentation. In our augmentation formulation, this can be achieved by increasing the augmentation mask ratio ($m$), which controls the fraction of features randomly replaced with features from other samples of the same class.
While we initially set a maximum value of $m=0.2$ during hyperparameter tuning, we further explore the impact of larger values of $m$ on the APIGraph and BODMAS datasets, which correspond to the Android and Windows feature sets, respectively. Specifically, we vary $m$ from $0.0$ to $1.0$ while keeping all other hyperparameters fixed at their optimal values for these datasets. Figure 14 presents the F1-scores on the validation set for both datasets. Notably, even for large values of $m$, performance does not degrade due to augmentation. For instance, when $m=1.0$, meaning that every feature in a sample is replaced with values drawn from other samples of the same class, the model still outperforms the case where $m=0.0$ (i.e., no augmentation).
Moreover, we observe that preserving label consistency—by excluding samples where model predictions differ between the original and augmented versions—tends to yield better performance than when such samples are included. This further supports the idea that our augmentation strategy maintains meaningful feature relationships without introducing harmful inconsistencies.
Appendix K Theoretical Analysis
Our self-training approach for malware detection under concept drift is theoretically supported by prior work on gradual self-training [39]. We outline key theoretical results explaining its effectiveness under distribution shifts, with commonly used notations summarized in Table XXII.
TABLE XXII: Summary of Notations
Notation Description
$P,Q$ Source and target data distributions, respectively
$P_{t}$ Data distribution at time-step $t$
$\theta,\theta^{\prime}$ Original and self-trained (updated) classifiers, respectively
$\Theta_{R}$ Hypothesis class constrained by regularization parameter $R$
$\rho(P,Q)$ Wasserstein distance between distributions $P$ and $Q$
$\rho_{m},\rho_{b}$ Wasserstein distances for malware and benign classes
$R$ Regularization parameter
$B$ Upper bound on the norm of feature vectors
$n$ Number of unlabeled samples from target distribution
$n_{\tau_{m},\tau_{b}}$ Number of samples retained after applying class-specific thresholds
$\alpha^{*}$ Minimum achievable ramp loss on the target domain
$\alpha_{0}$ Initial ramp loss on the source domain
$L_{r}(\theta,P)$ Ramp loss of classifier $\theta$ on distribution $P$
$L_{r}^{m}(\theta,P),\,L_{r}^{b}(\theta,P)$ Class-specific (malware and benign) ramp losses
$q_{m},q_{b}$ Proportions of malware and benign samples ($q_{m}+q_{b}=1$)
$\tau_{m},\tau_{b}$ Confidence thresholds for malware and benign predictions, respectively
$\beta$ Error amplification factor, defined as $\frac{2}{1-\rho R}$.
$\delta$ Confidence parameter for probabilistic bounds
K-A Setup and Assumptions
We consider a gradual domain shift setting where we have a sequence of distributions $P_{0},P_{1},\dots,P_{T}$ with $P_{0}$ as the source domain and $P_{T}$ as the target domain. The shift is assumed to be gradual, meaning that for a small $\epsilon>0$, the distance $\rho(P_{t},P_{t+1})<\epsilon$ for all $0\leq t<T$, where $\rho$ is a distributional distance metric.
We assume the existence of a sequence of classifiers $\theta_{t}$ that can classify most samples correctly with a margin and that the shift between consecutive domains is small enough to allow iterative adaptation.
K-B Main Theoretical Results
Theorem 1 ([39], Theorem 3.2)
$L_{r}(\theta^{\prime},Q)$ $\leq\frac{2}{1-\rho R}L_{r}(\theta,P)+\alpha^{*}$
$\quad+\frac{4BR+\sqrt{2\log(2/\delta)}}{\sqrt{n}},$ (6)
Corollary 1 ([39], Corollary 3.3)
K-C Extension to ADAPT
K-C1 Assymetric Error
Following Lemma A.2 from [39], we analyze the classification error for each class separately under these different shift magnitudes. The classification error on the target distribution $Q$ for malware and benign classes can be bounded as:
$\text{Err}^{m}(\theta,Q)$ $\leq\frac{1}{1-\rho_{m}R}L_{r}^{m}(\theta,P),$
$\text{Err}^{b}(\theta,Q)$ $\leq\frac{1}{1-\rho_{b}R}L_{r}^{b}(\theta,P).$ (8)
Since $\rho_{m}>\rho_{b}$, it follows that the multiplicative factor $\frac{1}{1-\rho_{m}R}$ is larger than $\frac{1}{1-\rho_{b}R}$. This implies that, even if the initial ramp losses $L_{r}^{m}(\theta,P)$ and $L_{r}^{b}(\theta,P)$ were similar, the resulting classification error for malware will be higher than that for benign samples under distribution shift.
A direct consequence of this is that malware samples, experiencing a greater shift, are more likely to be pushed toward the benign decision region than the reverse. Since errors occur primarily when samples cross the decision boundary, a higher shift for malware leads to more malware samples being misclassified as benign, increasing the false negative rate (FNR). In contrast, the smaller shift $\rho_{b}$ for benign samples means fewer of them cross into the malware region, resulting in a lower false positive rate (FPR). Thus, the asymmetry in shift magnitudes naturally leads to FNR exceeding FPR, explaining the observed trend in malware classification under distribution shift.
K-C2 Impact of Class-Specific Thresholding
$L_{r}^{m}(\theta,Q_{\tau_{m}}^{m})\leq L_{r}^{m}(\theta,Q),$
$L_{r}^{b}(\theta,Q_{\tau_{b}}^{b})\leq L_{r}^{b}(\theta,Q).$ (9)
K-C3 Modified Error Bound
First, the error term in Equation (6) becomes:
$\quad+q_{b}\frac{2}{1-\rho_{b}R}L_{r}^{b}(\theta,P_{\tau_{b}}).$ (12)
Thus, the final modified bound becomes:
$\quad+q_{b}\frac{2}{1-\rho_{b}R}L_{r}^{b}(\theta,P_{\tau_{b}})$
$\quad+\alpha^{*}$
$\quad+\frac{4BR+\sqrt{2\log(2/\delta)}}{\sqrt{n_{\tau_{m},\tau_{b}}}}.$ (13)
The modified bound in Equation (13) reduces to the original bound in Equation (6) when the distribution shifts for both classes are identical ($\rho_{m}=\rho_{b}=\rho$) and all samples are selected after thresholding, i.e., no data is filtered ($Q_{\tau_{m},\tau_{b}}=Q$, $n_{\tau_{m},\tau_{b}}=n$).
K-C4 Impact of Augmentation & Mixup
Data augmentation and mixup have been shown to act as regularizers during training [41, 76], effectively reducing overfitting. This regularization effect leads to a decrease in the bound given in Equation 13. Specifically, stronger regularization implicitly reduces the norm of the learned weight vector $w$, which corresponds to a smaller $R$ in our formulation (i.e., enforcing $\|w\|_{2}\leq R$). A smaller $R$ limits the model’s capacity, preventing it from fitting noise or learning overly complex decision boundaries. As a result, stronger regularization (smaller $R$) tightens the bound and enhances robustness to distribution shifts. We leave a precise theoretical analysis of these effects for future work.
Appendix L Hyperparameter Search
We present the range of hyperparameters used for different methods in Table XXIII. For self-training baselines, we adopt the default parameters for the five online learning classifiers used in our experiments. Insomnia, MORSE, and our proposed method ADAPT utilize different underlying machine learning models. Specifically, Insomnia employs a co-training approach with XGBoost and MLP, while MORSE uses MLP. In contrast, we evaluate ADAPT with Random Forest, XGBoost, and MLP. The same parameter search space is applied for these models as in the offline learning setup. We conduct a random search of 200 trials within the joint search space (i.e., combining classifier-specific and self-training-specific parameters) to select the best hyperparameters for all datasets.
TABLE XXIII: Hyperparameter search spaces for various methods. ”U” indicates sampling from a uniform random distribution within the specified range. Hyperparameters marked with ”*” are only used during the retraining phase. The Insomnia method also uses hyperparameters from MLP and XGBoost, while MORSE uses those from MLP. ADAPT utilizes the baseline model, which can be either Random Forest (RF), XGBoost, or MLP.

### v2

@@ X Limitations & Future Work
Evasion and Poisoning Attacks. The effectiveness of our approach may be compromised by adversarial samples designed to evade malware detection classifiers [37, 73]. While we do not explicitly address adversarial examples, integrating adversarial robustness techniques [13] could enhance resilience. Additionally, poisoning attacks, where an adversary injects carefully crafted samples into the unlabeled data pool, can impact continual learning updates under concept drift [67, 42]. Such poisoning can cause the pseudo-labeling process to reinforce incorrect predictions, increasing the risk of persistent false negatives or false positives. Although semi-supervised learning methods may offer some resistance to random label noise [76], defending against targeted adversarial poisoning remains an open challenge [34]. We leave the exploration of adversarial robustness and poisoning-aware defenses in pseudo-labeling pipelines as important directions for future work.
Computational Efficiency. ADAPT introduces no additional latency during deployment, as its inference time remains identical to that of the baseline model. However, the training time increases linearly due to data augmentation and mixup strategies. We provide a detailed analysis of the computational overhead in Appendix F.
@@
Catastrophic forgetting. Catastrophic forgetting refers to the tendency of a learned model to lose previously acquired knowledge when trained on new data—a challenge particularly relevant in malware detection [58]. We analyze its impact on the Android malware detection datasets in Appendix E. While our approach does not directly address catastrophic forgetting, future work could jointly integrate mechanisms to handle concept drift and mitigate forgetting. One promising direction is using replay-based methods from continual learning [59]. In our context, this could involve maintaining a replay buffer of class-balanced, confidently pseudo-labeled samples from intermediate time steps, which can be reused during subsequent model fine-tuning.
@@ XI Conclusion
We present ADAPT, a pseudo-labeling-based approach to mitigate concept drift in malware detection. Our model-agnostic framework is validated across five diverse datasets exhibiting different types and degrees of distributional shift. Through comprehensive experiments, we show that ADAPT overcomes limitations of prior pseudo-labeling strategies and consistently improves detection performance. Our method also integrates well with active learning, further enhancing adaptability in practice. Overall, ADAPT provides a practical, label-efficient, and robust solution for maintaining detection accuracy under evolving threats, and its principles are readily extensible to other security and drift-sensitive tasks.
Acknowledgment
The authors acknowledge Research Computing at the Rochester Institute of Technology [53] for providing computational resources and support that have contributed to the research results reported in this publication.
Appendix A Self-Training Baselines
DroidEvolver [77] and its improved version, DroidEvolver++ (DE++) [41], are notable self-training approaches for Android malware detection, but both suffer from self-poisoning and limited evaluation across datasets. Beyond these, research on self-training for concept drift in malware detection remains scarce, motivating our inclusion of additional established self-training algorithms for comparison.
@@ Appendix A Self-Training Baselines
    Adaptive Random Forest (ARF) [35]: ARF is an online ensemble method for concept drift adaptation. To enable self-training, we introduce a threshold parameter for pseudo-label selection, which is tuned in conjunction with other tree hyperparameters, allowing ARF to update using pseudo-labeled samples without requiring ground-truth labels.
@@
    DroidEvolver++ (DE++) [41]: DE++ improves upon DroidEvolver [77] for malware drift adaptation without requiring ongoing ground truth. It uses an ensemble of five linear models to generate pseudo-labels and tracks model aging with a buffer of recent samples. When prediction consistency in the buffer drops below a threshold, the model is updated using new pseudo-labels from the ensemble.
@@
    Insomnia [9]: Insomnia adapts to concept drift in network intrusion detection via co-training between a Nearest Centroid (NC) classifier and an MLP, with the NC providing pseudo-labels for uncertain MLP predictions. Both models are updated using pseudo-labeled data. For malware detection, we found NC to be ineffective, so we use XGBoost with the MLP, and select the most confident samples for updating to avoid early self-poisoning.
@@
    MORSE [76]: MORSE is a state-of-the-art semi-supervised method for malware family classification with noisy labels. It uses the FixMatch [65] framework, incorporating an augmentation strategy that enforces consistency between weakly and strongly augmented versions of each pseudo-labeled sample. We include MORSE as a baseline due to its strong performance in the presence of label noise.
The first two baselines, ARF and DE++, are online methods that update one sample at a time and do not require access to the full training dataset after initialization. In contrast, Insomnia and MORSE use batch training, updating models by combining original training data with pseudo-labeled samples. While our main experiments assume the training data remains available, we also consider restricted scenarios without access to the original training data (see Appendix D).
Appendix B Multiclass Classification
TABLE X: Statistics of the BODMAS multiclass classification dataset
ID Family
Train
Samples
Validation
Samples
Test
Samples
@@ 10 Berbew 284 195 1195
TABLE XI: Statistics of the EMBER multiclass classification dataset
ID Family
Train
Samples
Test
Samples
@@ 10 Tiggre 405 1284
B-A Dataset Statistics
Table X and Table XI present class-wise statistics for the BODMAS and EMBER datasets, respectively.
For BODMAS, we utilize all available samples per malware family in training and randomly sample 1,000 benign samples per month across all splits. Some families, like Gepys, are rare in the test set, while others, such as Berbew, are more frequent during testing than training. For EMBER, we cap each class, including Benign, at 500 samples in training and per month in testing to manage dataset size. This produces approximately balanced training data but class imbalance in the test set.
B-B Experiments Excluding Benign Class
In practice, malware analysis typically proceeds in two steps: first, binary classification distinguishes between benign and malicious samples; second, if a sample is identified as malware, its family is determined [58]. We emulate this scenario for malware family classification by removing benign samples from the BODMAS (Table X) and EMBER (Table XI) datasets, resulting in a 10-class task. In this setting, ADAPT uses a single threshold parameter $\tau_{m}$ shared across all malware families; otherwise, the algorithm is unchanged from the binary case. This reduces the number of ADAPT hyperparameters from five to four.
TABLE XII: Performance on multiclass classification (10-class) on BODMAS and EMBER datasets.
@@ B-B Experiments Excluding Benign Class
TABLE XIII: Performance on multiclass classification (10-class) on BODMAS and EMBER datasets with active learning.
@@
Table XII presents the results of ADAPT and other self-training approaches across both datasets. ADAPT outperforms the baseline XGBoost methods on both datasets, while other self-training methods perform worse. The improvement is especially notable on the EMBER dataset, where ADAPT achieves a 6.5% increase in the F1-score. Table XIII presents the results for active learning with a monthly annotation budget of 50 samples. Active learning with uncertainty sampling outperforms the baseline XGBoost model trained with the same annotation budget using random sampling. Additionally, incorporating ADAPT further enhances performance on both datasets. Specifically, integrating ADAPT improves the F1-score by 2.3% on the BODMAS dataset and 6.1% on the EMBER dataset, demonstrating the effectiveness of leveraging unlabeled samples even in the presence of active learning.
Appendix C Support Vector Machine(SVM) Experiments
We conduct additional experiments using SVM baseline models on two Android datasets, APIGraph and Drebin, using the same hyperparameter selection strategy from Section VI-C. The results, including those with ADAPT, are summarized in Table XIV.
The baseline SVM model performs competitively with the baseline XGBoost model on the APIGraph dataset. However, its performance is significantly lower on the Drebin dataset than the XGBoost baseline.
@@ Appendix C Support Vector Machine(SVM) Experiments
TABLE XIV: SVM performance on APIGraph and Drebin (average over five runs).
APIGraph Drebin
Method F1 FPR FNR F1 FPR FNR
SVM 66.3 1.41 39.5 33.8 0.70 76.7
SVM+ADAPT 73.6 1.47 29.3 48.0 0.84 61.5
Appendix D Source-Free Adaptation
@@
In some scenarios, access to the original training data may be unavailable after initial model training due to privacy, proprietary, storage, or computational constraints [77]. In these cases, model updates must rely solely on pseudo-labeled data. We explore this setting for XGBoost on the Drebin and APIGraph datasets, modifying Algorithm 1 at line 7 so that $D_{m}=D_{p}$, meaning only pseudo-labeled data are available after the initial phase. Separate hyperparameter optimization is performed for this restricted scenario using the same search space.
Figure 13 compares the baseline XGBoost model, the ADAPT semi-supervised update, and source-free adaptation on these datasets. For APIGraph, results are shown for the first 18 test months, as performance declines over longer periods (average F1: 61.24%), though the source-free method offers noticeable gains in this window. On Drebin, source-free adaptation improves the average F1 by 11.3% over the baseline in the first year, lagging the semi-supervised variant by just 1.76%. On APIGraph, it outperforms the baseline by 0.82% but is 3.97% below the semi-supervised approach. These findings indicate that ADAPT can still mitigate concept drift when the original training data is unavailable, making it suitable for continual learning scenarios with storage limitations [58].
@@
Figure 14: Impact of catastrophic forgetting on Drebin: XGBoost and MLP performance over the first 12 test months. $M_{T}$ is the final-month model; $M_{i-1}$ is the previous month’s model.
@@
Appendix E Catastrophic Forgetting
Catastrophic forgetting, where a model loses previously acquired knowledge when learning from new data, is a significant challenge in continual learning [58]. While our work does not directly address this issue, we assess its impact on the Android malware datasets using XGBoost and MLP. Specifically, we evaluate the final model trained on the last month ($M_{T}$) against the first year of test data, comparing its F1 score, FPR, and FNR to those of models trained during each monthly retraining phase.
@@ Appendix E Catastrophic Forgetting
Appendix F Computational Efficiency
The extent of this increase depends on the dataset size, feature dimensionality, and model complexity. Table XV reports the training time overhead introduced by ADAPT across different models and datasets. The most significant increase is observed for the XGBoost model on the Drebin dataset, where the high feature dimensionality (16,978) substantially impacts computation time. For other datasets, training time typically increases by three to five relative to the baseline model.
TABLE XV: Training time (in seconds) of different baseline models with ADAPT
Dataset RF
RF +
ADAPT
MLP
MLP +
ADAPT
XGBoost
XGBoost +
ADAPT
@@ Appendix F Computational Efficiency
Appendix G Hyperparameter Search
Table XVI summarizes the hyperparameter ranges used for each method. For self-training baselines, we adopt default parameters for the five online learning classifiers. Insomnia, MORSE, and our method ADAPT use different underlying models: Insomnia applies co-training with XGBoost and MLP, MORSE uses MLP, and ADAPT is evaluated with Random Forest, XGBoost, and MLP. All models share the same parameter search space as in the offline setting. We perform a random search with 200 trials across the joint parameter space to select the best hyperparameters for each dataset.
TABLE XVI: Hyperparameter search spaces for various methods. ”U” indicates sampling from a uniform random distribution within the specified range. Hyperparameters marked with ”*” are only used during the retraining phase. The Insomnia method also uses hyperparameters from MLP and XGBoost, while MORSE uses those from MLP. ADAPT utilizes the baseline model, which can be either Random Forest (RF), XGBoost, or MLP.